OpenStudy (anonymous):

Superman needs to save Lois from the clutches of Lex Luthor. After flying for 6 seconds, he is 1900 meters from her. Then at 13 seconds he is 1550 meters from her.

OpenStudy (mathmale):

JANB: Are you sure that this is all the info supplied? I don't see any question here!

OpenStudy (anonymous):

Write an equation to model this situation (use m for meters and s for seconds).

OpenStudy (mathmale):

You are given 2 sets of info: (6 seconds, 1900 meters) and (13 seconds, 1550 meters). You may as well assume that this guy flies in a straight line (even though i don't). So you are given two points; you need to imagine that our hero flies from one point to the next. You need to come up with a mathematical model that describes his inflight adventure. How do you describe a straight line?

OpenStudy (anonymous):

Its 180 degrees? I am honestly really lost with this question.

OpenStudy (mathmale):

if you were to fly from Los Angeles to New York, your flight path would have "slope," since NY is north of, as well as to the east of, LA. How would you find that slope? Let me estimate that NY is 800 miles north of LA and 3000 miles east of LA. What is the slope of that flight line?

OpenStudy (anonymous):

That's extremely confusing to be honest. y=800x+3000? no? I don't know?

OpenStudy (anonymous):

Calculate the slope first.\[\frac{ 1900 -1550 }{ 13-6 }=50\]This means Superman is flying at a rate of 50 m/s. Then you use the point-slope formula by picking one of the two points.\[y-1550=50(x-6)\] I'll leave it to you to convert this in slope-intercept form.

OpenStudy (anonymous):

I got 50 but I was not sure what to do next. I should've known to use the y-y1=m(x-x1) I'm such an idiot !

OpenStudy (mathmale):

In fairness, I should present (or review) the concept of slope. It helps to think of "slope" as rise over run. To get to NY from LA, you'd have to fly 800 miles north and at the same time 3000 miles east. The "rise" (motion in the vertical direction) is 800 miles; the "run" (motion in the horizontal direction) is 3000 miles. Therefore, the slope, m, is \[m=\frac{ rise }{ run }=\frac{ 800 miles }{ 3000 miles }=\frac{ 8 }{ 30 }=\frac{ 4 }{ 15 }.\]

OpenStudy (mathmale):

Have you seen this formula before?

OpenStudy (anonymous):

Yeah I have, I just am not really sure when to use what formula if that makes any sense.

OpenStudy (mathmale):

In the math problem you want to solve: your initial point is (6,1900); your final point is (13,1550). Here the 6 and 13 represent TIME and are on the horizontal axis; the 1900 and the 1550 represent how far SuperGuy is from Lois. The "run" is the difference between 6 and 13: 13-6=7. The "rise" is the difference between 1900 and 1550: 1550-1900=350./ thus, the slope is m = rise/run, or 7/(-350), or -7/350.

OpenStudy (mathmale):

I must apologize: I have the order wrong in my statement, "the slope is -7/350." I should have typed, the slope is -350 meters / 7 seconds, or -50 meters/second. sorry.

OpenStudy (triciaal):

negative distance with positive time the slope should be negative as he gets closer the time is increasing

OpenStudy (mathmale):

How are you doing? Does this make sense, or do you need clarification?

OpenStudy (mathmale):

tricia: Agreed. The slope is -50 meters/ sec.

OpenStudy (mathmale):

As Tricia says, SuperGuy's distance from the hapless Lois decreases with time, at the rate of 50 meters/sec.

OpenStudy (mathmale):

Jean: Awaiting a response from you.

OpenStudy (anonymous):

I used the information you gave me and came up with this y-1550=-50(x-13) ?

OpenStudy (anonymous):

Am I using it wrong ?

OpenStudy (triciaal):

the objective i think is to find x when y = 0 how long will it take to rescue Lois to get from

OpenStudy (triciaal):

you have 2 unknowns in what you have written

OpenStudy (triciaal):

this is what I would normally do after finding the slope use one of the given points to find the intercept using the equation of the line find other points on the line slope = -50 using (6, 1900) 1900 = -50(6) + b 1900 + 300 = b = 2200 y = -50(x) + 2200

OpenStudy (triciaal):

the question asked only for the equation

OpenStudy (mathmale):

I agree: "Write an equation to model this situation (use m for meters and s for seconds)."

OpenStudy (anonymous):

I have been on this problem for hours trying to figure out what I needed to do. But I think I get it now. You plugged in what you found into the slope intercept form, that makes much more sense.

OpenStudy (triciaal):

distance traveled in meters,m = 50s + 2200

OpenStudy (mathmale):

Tricia, wouldn't the independent variable be time (t)? Yes, we're measuring time in seconds (s), but nevertheless the equation should be d = distance = [(-50 meters)/sec]t + 2200 meters.

OpenStudy (triciaal):

from the information given , the 2 points, find the slope of the line we have points and the slope we can find the equation of the line

OpenStudy (mathmale):

I'm not arguing with your result: distance traveled in meters,m = 50s + 2200 But I am stating that units of measurement are required, and that your equation would be better if written as \[d=(\frac{ -50 meters }{ \sec })t + 2200 meters,\] where t is, of course, measured in seconds (s).

OpenStudy (anonymous):

Thank you, both of you by the way !

OpenStudy (triciaal):

@mathmale yes that would be more precise. @jeaannebad your welcome